OpenStudy (anonymous):

A 5.00-grain aspirin tablet has a mass of 320 mg. For how many kilometers would the energy equivalent of this mass power an automobile? Assume 12.75 km/L and a heat of combustion of 3.65x10^7 J/L for the gasoline used in the automobile.

OpenStudy (agent0smith):

Convert the mass (320mg = 0.32kg) to energy in joules J \[E=mc ^{2}\] The engine uses 3.65x10^7 J/L so we need to find how many litres of gasoline, G, that energy is equivalent to: \[\frac{ E }{3.65x10^7 J/L } = G\] Then we multiply the number of litres of gasoline, G, by the fuel efficiency 12.75 km/L to find the distance in kilometres that it'll take us: \[G \times 12.75 km/L = distance\]

OpenStudy (agent0smith):

oops, that mass should be 320mg = 0.32g = 0.00032 kg

OpenStudy (anonymous):

Thanks.